OpenStudy (anonymous):

Tell me if I did this right Solve using the substitution method 6x+5y=-26 x=-2-2y the solution is (2,-6)

OpenStudy (anonymous):

6(-2-2y)+5y=-26 -12-12y+5y=-26 -12-7y=-26 -7y=-14 y=2 x=-2-2(2) x=-2-4 x=-6

OpenStudy (anonymous):

nope the solution is (-6,2)

OpenStudy (anonymous):

yeah just saw the numbers ;S

OpenStudy (anonymous):

That is how you solve using the substitution method.